OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 4, -8, and 2 + 3i

OpenStudy (anonymous):

\[(x-4)(x+8)(x-(2+3i))(x-(2-3i))\] is a start

OpenStudy (anonymous):

first product is easy enough, you get \[(x-4)(x+8)=x^2-4x-32\] it is the second that looks difficult, but it isn't

OpenStudy (anonymous):

for the other 2 how do you work that out?

OpenStudy (anonymous):

\[(x-(2+3i))(x-(2-3i))\] use first \(x^2\) last \(2+3i)(2-3i)=2^2+3^2=4+9=13\) outer and inner you will get \((-2+3i)x+(-2-3i)x\) and you see you are left with only \(-4x\)

OpenStudy (anonymous):

don't forget that when you multiply a complex number \(a+bi\) by its complex conjugate \(a-bi\) you get \(a^2+b^2\) a real number, so last term is easy, you can do it in your head

OpenStudy (anonymous):

then you can see that \(-(a+bi)x+-(a-bi)x=-2ax\)

OpenStudy (anonymous):

woah you lost me

OpenStudy (anonymous):

so last two terms give \[x^2-4x+13\]

OpenStudy (anonymous):

oh ok the way i was reading it was confusing

OpenStudy (anonymous):

you know you have to do this multiplication \((x-(2+3i))(x-(2-3i))\) right?

OpenStudy (anonymous):

yea the product of that gives you -4x

OpenStudy (anonymous):

i hate to say "foil" but that is the idea, first outer inner last first is \(x^2\) last is \(2^2+3^2\) and outer plus inner is \(-4x\)

OpenStudy (anonymous):

finally you get to multiply \[(x^2-4x-32)(x^2-4x+14)\] to finish

OpenStudy (anonymous):

yea i undertand i was just reading it weird

OpenStudy (anonymous):

ok have fun

OpenStudy (anonymous):

then you multiply (x2 - 4x +32)(x2-4x+13) right?

OpenStudy (anonymous):

how do you multiply them?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Good ol' distributive property, just go term by term and group like terms together after multiplying.

OpenStudy (anonymous):

It's no different than multiplying 123 by 456 by expressing the product as (100 + 20 + 3) × (400 + 50 + 6) =100×400 + 100×50 + 100×6 + 20×400 + 20×50, etc.